OpenStudy (anonymous):

Rust, chemically, is iron(III) oxide,Fe2O3.On a large ship, 16kg of rust accumulates in a day. Calculate how many kilograms of iron and how many kilograms of oxygen would be used up.

OpenStudy (anonymous):

Fe=4.19 O=.9024

OpenStudy (anonymous):

It's wrong.

OpenStudy (anonymous):

hint: the molar mass of Fe2O3 is 159.70 So, the number of mol of Fe2O3 is: n=16000/159.70= 100.18

OpenStudy (anonymous):

now, 1 mol of Fe2O3 contains 2 moles of Fe and 3 moles of O, so we can write: mFe=(100.18*2*55.85)/1000=11.2 mO=(100.18*3*16)/1000=4.8

OpenStudy (anonymous):

where mO and mFe are the masses of oxygen and iron respectively

OpenStudy (anonymous):

that's right! As we can see if we compute this sum: 11.2+4.8, we get the initial mass, namely 16 Kg